# IQA-Eval: Automatic Evaluation of Human-Model Interactive Question Answering

### v1

@@ Abstract
To evaluate Large Language Models (LLMs) for question answering (QA), traditional methods typically focus on directly assessing the immediate responses generated by the models based on the given question and context. In the common use case of humans seeking AI assistant’s help in finding information, these non-interactive evaluations do not account for the dynamic nature of human-model conversations, and interaction-aware evaluations have shown that accurate QA models are preferred by humans 24. Recent works in human-computer interaction (HCI) have employed human evaluators to conduct interactions and evaluations, but they are often prohibitively expensive and time-consuming to scale. In this work, we introduce an automatic evaluation framework IQA-Eval to Interactive Question Answering Evaluations [^1] https://github.com/du-nlp-lab/IQA-Eval, more specifically, we introduce LLM-based Evaluation Agent (LEA) that can: (1) simulate human behaviors to generate interactions with IQA models; (2) automatically evaluate the generated interactions. Moreover, we propose assigning personas to LEAs to better simulate groups of real human evaluators. We show that: (1) our evaluation framework with GPT-4 (or Claude) as the backbone model achieves a high correlation with human evaluations on the IQA task; (2) assigning personas to LEA to better represent the crowd further significantly improves correlations. Finally, we use our automatic metric to evaluate five recent representative LLMs with over 1000 questions from complex and ambiguous question answering tasks, which comes with a substantial cost of \$5k if evaluated by humans.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of natural language processing (NLP), enabling systems to perform a wide range of tasks with remarkable proficiency 55, 48, 5. Among these tasks, question answering (QA) has emerged as a critical and representative goal-oriented application, demonstrating the potential of LLMs to generate informative responses as an assistant 4. The success of these applications or in general the success of LLMs, heavily relies on accurate and consistent evaluation 8. Accurate and consistent evaluation helps researchers better understand LLMs capacities, and offers insights for new human-LLMs QA interactions 29, 8.
@@
Traditionally, automatic metrics such as accuracy have been used to evaluate models based on the quality of their direct answers to specific questions. However, as interactions between humans and LLMs grow more complex and nuanced, these traditional metrics often fail to capture the full spectrum of a model’s capabilities (e.g. helpfulness and providing enjoyment), particularly in interactive QA settings 30, 12, whereas interactions are crucial for user experience and system effectiveness, yet remains overlooked in traditional evaluation paradigms. Recently work such as 24 evaluated human-LLM interactions, a process that involves human participation and annotation. Although human evaluations for these interactions provide a closer approximation to real-world use cases, this approach is significantly costly and time-consuming and could also be prone to errors. Recognizing the need for automatic evaluation, works like G-Eval 32, LLM-Eval 29 and PRD 27 proposed to automate the assessment of non-interactive LLM responses using LLMs.
Drawing insights from (a) using LLM for automatic evaluation and (b) literature of LLM-agents research 46, 13, we propose IQA-Eval framework to auto-evaluate the performance of IQA models (i.e. LLMs) with LLM-based Evaluation Agent (LEA) to simulate and then evaluate interactions. By incorporating personas, our experiments on a well-annotated dataset show that our methods align well with human judgments and provide a more comprehensive evaluation of LLMs in interactive settings than traditional metrics. See an overview of our system in Figure 1.
Meta-evaluation shows that IQA-Eval results highly correlate with human evaluations across metrics and in general (with scores being over 0.6). The well-alignment with human judgments underscores IQA-Eval as a significant advancement in the field of automated IQA evaluation, offering a robust alternative to traditional/manual methods. Finally, we benchmark recent LLMs with new complex and ambiguous questions, and find that the accuracy of answers does not transfer to the corresponding ranking of models on their capability of achieving good human-model interactions.
@@
  • We propose assigning personas to the LLM evaluation agent, enabling a more nuanced and accurate assessment of a model’s ability to adapt its responses to real-world interaction scenarios, where the context and individual user preference significantly influence the nature of the interactions.
@@ 2 Related Work
Traditional methods for human-model dialogue evaluation have often been centered around single-turn pairwise evaluation 44, 25. Some methods with multi-turn Likert scores emerge 43, 54, 40 but require time-consuming and costly collection of human-model conversations. In response, 16 suggests a self-play scenario where a dialog system engages in conversation with itself, employing a set of proxy measures for evaluation.
Acute-eval 26 and LLM-Eval 29 have advanced multi-turn evaluation frameworks, reflecting the increasing demand for sophisticated techniques that holistically capture human-AI interactions. Further studies in specific interaction domains like task completion 31, code generation 52, and collaborative problem-solving 24, 19, 14 emphasize the need for evaluations that consider both environmental and human elements 47. Our work differs from these methods by introducing an automated approach that emphasizes interaction quality and significantly reduces the reliance on human annotations.
@@
Personas are constructed profiles that represent key traits of a group of users, as defined in the HCI field, reflecting their characteristics, behaviors, and goals to guide the design of technologies that are well-suited to user needs 11. This approach is a promising direction for enhancing relevance and personalization in NLP applications 36, 2, 41, 57, offering significant potential for customizing engagement and improving the effectiveness of conversational agents 25, 54, wolf2019transfertransfo, 7, 33, 56. Li et al. 25 introduces a persona-based neural conversation model to enhance dialogue personalization and coherence. Zhang et al. 54 develops personalized dialogue agents that incorporate user-specific details to enhance interaction. In our work, persona settings enable our framework to tailor interactions and assessments, aligning more closely with the specific characteristics and preferences of different user groups.
LLM-based Agent for Simulation
Recently, LLMs demonstrated human-like intelligence, as evidenced in various studies 51, 39, 28, 21, 3, 6, 42, 10. The integration of LLMs into agents that simulate complex human behaviors and social interactions is an area of growing research interest 34, 49, 50. For example, Park et al. 37 and Gao et al. 15 employ these agents to simulate human emotions, attitudes, and behaviors in social networks, while Park et al. 38 leverages in-context learning to simulate human behaviors in sandbox world. Moreover, Argyle et al. 1 utilizes ”algorithmic bias” inherent in GPT-3 to reflect the response patterns of different human subgroups. Horton 18 utilize LLMs in experimental setups of behavioral economics experiments to facilitate pilot studies. Additionally, Hämäläinen et al. 20 and Wang et al. 45 investigate LLM-based agents in recommender systems to simulate and collect data on user behavior. These studies show the broad applicability and potential of LLM-based agents in simulating human behaviors and interactions across diverse applications. Our work utilizes LLM-based evaluation agents(LEAs) to fully automate interactive quality assessments, handling both interaction generation and evaluation, to enhance the evaluation of IQA models in realistic scenarios.
@@ 3 IQA-Eval: Evaluating Interactive Question Answering (IQA)
In this section, we introduce our IQA-Eval framework for automatically evaluating Interaction Question Answering Models (IQA models) with LLM-based Evaluation Agents (LEA). IQA-Eval utilizes LEA to simulate humans in the following two stages: (1) generating interactions with IQA models; and (2) evaluating interactions. Additionally, we explore assigning personas to the IQA-Eval to better simulate the diversity of the groups of humans and provide individualized evaluations.
@@ 3.1 Interaction Generation with LEA (Stage 1)
Inspired by peer discussions 24, 46, we propose utilizing LEAs with prompt-based techniques to simulate human behaviors for effective interaction generation with IQA models. The structured prompt includes three key components: (1) a role description; (2) a task description; and (3) instructions for the discussion.
@@
Discussion instruction guides LEAs on their subsequent steps by providing detailed descriptions to facilitate progress in interactions. It comprises two essential components: (1) the actions to take; and (2) the detailed procedures to follow. For example, in a question answering task, the prompt specifies:
You can ask an assistant questions for help. Please ask sub-questions
to approach answers. In each turn, please only ask one sub-question
to interact with the assistant.
In the sub-question, please include all necessary information in the
original question, such as the question and all options. If you know
the answer, please output "So, the answer is: A, B, C, or D".
The first paragraph of the discussion instruction outlines “the actions to take” and the second paragraph details “procedure to follow”.
At the start of an interaction, the LEA receives a system prompt that includes all three components above, along with the specific question to be addressed. As the LEA interacts with the IQA model, it generates sub-questions to request clarification of unknown entities, definitions, or particular aspects of the original question. Then, the IQA Model takes the questions as the input and output responses. After receiving responses, the LEA continues to pose further questions until it determines the final answer. The full prompt structure and interaction details are provided in Appendix B.2.
@@ 3.2 Interaction Evaluation with LEA (Stage 2)
Inspired by G-eval 32, which demonstrates that evaluations by GPT models align closely with human assessments in NLG tasks, we propose utilizing LEAs for interaction evaluation. This LEA assesses interactions generated by LEAs and IQAs in Stage 1. The module takes task details, such as questions or articles, and interactions as the input, and output evaluation scores. The prompt contains three parts: (1) role and task description; (2) metrics definition; and (3) evaluation instruction.
@@
Metrics definitions describes the criterion that LEA needs to follow in the evaluation process. They can customized for different tasks. For a question answering task, we add the following prompt to define the “helpfulness” metric:
@@
Finally, all evaluation scores for metrics are calculated by averaging the results of multiple runs. The complete prompt is available in Appendix B.1, and further details about our implementation can be found in Section 4.
@@ 3.3 Assigning the Personas to LEA
The evaluation process for LLM IQA models, including interaction generation and interaction evaluation, typically utilizes a standard persona, designed to reflect neutral and average characteristics across various attributes such as knowledge, language proficiency, and beliefs. This approach ensures a consistent and uniform baseline for assessment. However, individual users often exhibit unique personal preferences and characteristics, making a one-size-fits-all evaluation less effective. Moreover, the persona distribution of the target user group significantly impacts the performance of IQA models in real-world applications. For instance, if 20% of human users prefer brief interactions and 70% prefer detailed information, applying a general LEA to simulate this group of persons is likely to result in poor correlation in terms of evaluations.
@@
For example, when the LEA is assigned with the “Critical-Seeker” persona (definition in B.3) for interaction generation, we adapt the default role and task description (in Stage 1) to:
You prefer interactions rich in critical information. You need help from an assistant and try to get critical information from it to answer the following questions. For interaction evaluation, the default role and task description prompt changes to The AI Assistant should provide straightforward, simple, and concise answers to aid users in deducing solutions. Additionally, the definition of metrics is also adjusted to align with this persona, with further details available in Appendix B.3.
@@ 4.1 Experiment Settings
We apply our evaluation method on the annotated dataset from the study by Lee et al. 24. This dataset consists of 3641 interactions from 331 annotators. Questions in the dataset are multi-choice and are derived from the MMLU dataset 17 (example question in Figure 1). The construction of MMLU requires each worker to participate in 11 random conversations with one of the following three IQA models: TextDavinci (text-davinci-001), TextBabbage (text-babbage-001), and Davinci (davinci-001). At the end of conversations, fluency and helpfulness scores are annotated by annotators. The number of queries and accuracy for each IQA model can be easily deduced from annotations. In this work, We adjust the four metrics to evaluate generated interactions:
@@
To evaluate the effectiveness of IQA-Eval framework, we experiment with different LEA models on the above-mentioned three LLM IQA models in MMLU. For LEA that conducts both interaction generation and interaction evaluation, we use ChatGPT (GPT-3.5-turbo-1106), GPT4 (GPT-4-1106-preview), Claude (Claude-1).
@@
To evaluate the performance of our IQA-Eval framework, we use Pearson correlations as the indicator between the Human evaluations and LEA evaluations of IQA models.
@@ 4.2 Experiment Results
According to Table 2, all models, including GPT4, GPT3.5, and Claude, show high correlation with human evaluations. GPT4 aligns most closely with human judgments in both “Helpfulness” and “Fluency” metrics and the highest overall correlation score. This indicates that these models are capable of effectively performing IQA-Eval framework as LEA models.
@@
In Table 1, ChatGPT scores closest to human judgments, particularly in the ”Helpfulness” metric. Conversely, GPT4 and Claude score lower on ”Helpfulness” than human evaluations, because they tend to produce inaccurate and repetitive responses that lack coherence and do not directly address user queries, as indicated by their generated explanations. IQA-Eval scores on “Fluency” are close and highly correlated to human judgments. Both scores given by human and LEA models show that IQA models provide fluent outputs. Furthermore, according to the ”# Queries” metric, most models conclude conversations more quickly than humans, except for Claude, which requires more turns, potentially due to its non-OpenAI origins that it needs more turns to adapt the conversational style and understand responses. Notably, GPT4 achieves the highest accuracy among all models.
@@
Results are in shown Figures 2(a) and 2(b). The Pearson correlation coefficients for fluency and helpfulness between human judgments and LEA evaluations show distinct patterns. ChatGPT demonstrates a stronger correlation with human ratings, recording a correlation score of 0.424 for fluency and a correlation score of 0.306 for helpfulness. In contrast, Claude shows slightly lower correlations (0.281 for fluency and 0.287 for helpfulness. These results indicate that ChatGPT aligns better with human evaluative standards compared to Claude in these specific metrics.
Moreover, for interactions between humans and IQA models (Figure 2), LEA evaluations moderately correlate with human evaluations. However, for interactions between LEAs and IQA models (Table 2), which is the main focus of our paper, LEA evaluations highly correlate (around 0.6) with human evaluations. This indicates that LEA models are helpful when participating in the whole evaluation process, including both the interaction and evaluation. However, when evaluating interactions between human and IQA models, LEAs focus differently from humans, which causes moderate correlations.

### v2

@@ Abstract
To evaluate Large Language Models (LLMs) for question answering (QA), traditional methods typically focus on assessing single-turn responses to given questions. However, this approach doesn’t capture the dynamic nature of human-AI interactions, where humans actively seek information through conversation [^1] Details are in Appendix F. Recent works in human-computer interaction (HCI) have employed human evaluators to conduct interactions and evaluations, but they are often prohibitively expensive and time-consuming to scale. We introduce an automatic evaluation framework IQA-Eval to achieve Interactive Question Answering Evaluations [^2] https://github.com/du-nlp-lab/IQA-Eval, more specifically, we introduce a LLM-based Evaluation Agent (LEA) that can: (1) simulate human behaviors to generate interactions with IQA models; (2) automatically evaluate the generated interactions. Moreover, we propose assigning personas to LEAs to better simulate groups of real human evaluators. We show that: (1) our evaluation framework with GPT-4 (or Claude) as the backbone model achieves a high correlation with human evaluations on the IQA task; (2) assigning personas to LEA to better represent the crowd further significantly improves correlations. Finally, we use our automatic metric to evaluate five recent representative LLMs with over 1000 questions from complex and ambiguous question answering tasks, which comes with a substantial cost of \$5k if evaluated by humans.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of natural language processing (NLP), enabling systems to perform a wide range of tasks with remarkable proficiency 65, 56, 63, 14, 26. Among these tasks, question answering (QA) has emerged as a critical and representative goal-oriented application, demonstrating the potential of LLMs to generate informative responses as an assistant 4. Multiple methods have been proposed to enhance the faithfulness and explainability of generated information 57, 39, 32, 13. Beyond developing these methods, rigorous evaluation of the generated outputs is also crucial.
@@
Accurate and consistent evaluation helps researchers understand existing LLM’s capacities and emerging human-LLM QA interactions 7, 35, 7. Traditionally, automatic metrics such as accuracy have been used to evaluate models based on the quality of their direct answers to specific questions. However, as interactions between humans and LLMs grow more complex and nuanced, these traditional metrics often fail to capture the full spectrum of a model’s capabilities (e.g. helpfulness and fluency), particularly in interactive QA settings 36, 11, whereas interactions are crucial for user experience and system effectiveness, yet remains overlooked in traditional evaluation paradigms. Recent works such as 29 evaluated human-LLM interactions, a process that involves human participation and annotation. Although human evaluations for these interactions provide a closer approximation to real-world use cases, this approach is significantly costly and time-consuming. Recognizing the need for automatic evaluation, works like G-Eval 38, LLM-Eval 35, FaithScore 25, and PRD 33 proposed to automate the assessment of non-interactive LLM responses using LLMs as evaluators.
Drawing insights from (a) using LLM for automatic evaluation and (b) literature of LLM-agents research 54, 12, we propose IQA-Eval framework to auto-evaluate the performance of IQA models (i.e. LLMs) with LLM-based Evaluation Agent (LEA) to simulate and then evaluate interactions. By additionally incorporating personas, our experiments on a well-annotated dataset show that our methods align well with human judgments and provide a more comprehensive evaluation of LLMs in interactive settings than traditional metrics. Finally, we benchmark recent LLMs with new complex and ambiguous questions, and demonstrate that the accuracy of answers does not always transfer to the corresponding ranking of models on their capability of achieving good human-model interactions. See an overview of our system in Figure 1.
@@
  • We propose persona-based LLM evaluation agents to better assess how models adapt to different user preferences and interaction styles.
@@ 2 Related Work
Traditional methods for human-model dialogue evaluation have often been centered around single-turn pairwise evaluation 52, 30. Some methods with multi-turn Likert scores emerge 51, 64, 48 but require time-consuming and costly collection of human-model conversations. In response, 19 suggests a self-play scenario where a dialog system engages in conversation with itself, employing a set of proxy measures for evaluation.
Acute-eval 31 and LLM-Eval 35 have advanced multi-turn evaluation frameworks, reflecting the increasing demand for sophisticated techniques that holistically capture human-AI interactions. Further studies in specific interaction domains like task completion 37, code generation 61, and collaborative problem-solving 29, 22, 15 emphasize the need for evaluations that consider both environmental and human elements 55. Our work differs from these methods by introducing an automated approach that emphasizes interaction quality and significantly reduces the reliance on human annotations.
LLM-based Agent for Simulation
Recently, LLMs rises to demonstrate human-like intelligence, as evidenced in various studies 60, 47, 34, 24, 3, 5, 50, 9. The integration of LLMs into agents that simulate complex human behaviors and social interactions is an area of growing research interest 41, 58, 59. For example, Park et al. 45 and Gao et al. 17 employ these agents to simulate human emotions, attitudes, and behaviors in social networks, while Park et al. 46 leverages in-context learning to simulate human behaviors in sandbox world. Moreover, Argyle et al. 1 utilizes “algorithmic bias" inherent in GPT-3 to reflect the response patterns of different human subgroups. Horton 21 utilize LLMs in experimental setups of behavioral economics experiments to facilitate pilot studies. Additionally, Hämäläinen et al. 23 and Wang et al. 53 investigate LLM-based agents in recommender systems to simulate and collect data on user behavior. These studies show the broad applicability and potential of LLM-based agents in simulating human behaviors and interactions across diverse applications. Our work utilizes LLM-based evaluation agents (LEAs) to fully automate interactive quality assessments, handling both interaction generation and evaluation, to enhance the evaluation of IQA models in realistic scenarios.
@@
Personas are constructed profile prompts that represent key traits of a group of users, as defined in the HCI field, reflecting their characteristics, behaviors, and goals to guide the design of technologies that are well-suited to user needs 10. This approach enhances relevance and personalization in NLP applications 43, 2, 49, 68, offering significant potential for customizing engagement and improving the effectiveness of conversational agents 30, 64, 6, 40, 67. Li et al. 30 introduces a persona-based neural conversation model to enhance dialogue personalization and coherence. Zhang et al. 64 develops personalized dialogue agents that incorporate user-specific details to enhance interaction. In our work, persona settings enable our framework to tailor interactions and assessments, aligning more closely with the specific characteristics and preferences of different user groups.
@@ 3 IQA-Eval: Evaluating Interactive Question Answering (IQA)
In this section, we introduce our IQA-Eval framework for automatically evaluating Interaction Question Answering Models (IQA models) with LLM-based Evaluation Agents (LEA). LEAs are used to simulate humans in the following two stages: (1) generating interactions with IQA models; and (2) evaluating interactions. Lastly, we discuss the use of personas to for LEAs.
@@ 3.1 Interaction Generation with LEA (Stage 1)
Inspired by peer discussions 29, 54, we prompt LEAs to simulate human behaviors for effective interaction generation with IQA models. The structured prompt includes three key components: (1) a role description; (2) a task description; and (3) instructions for the discussion.
@@
Discussion instruction guides LEAs on their subsequent steps by providing detailed descriptions to facilitate progress in interactions. It comprises two essential components: (1) the actions to take; and (2) the detailed procedures to follow. For example, in a question answering task, the prompt specifies: You can ask an assistant questions for help. Please ask sub-questions to approach answers. In each turn, please only ask one sub-question to interact with the assistant.
In the sub-question, please include all necessary information in the original question, such as the question and all options. If you know the answer, please output "So, the answer is: A, B, C, or D".
At the start of an interaction, the LEA receives a system prompt that includes all three components above, along with the specific question to be addressed. As the LEA interacts with the IQA model, it generates sub-questions to request clarification of unknown entities, definitions, or particular aspects of the original question. Then, the IQA Model takes the questions as the input and output responses. After receiving responses, the LEA continues to pose further questions until it determines the final answer. The full prompt structure and interaction details are provided in Appendix C.2.
@@ 3.2 Interaction Evaluation with LEA (Stage 2)
Inspired by G-eval 38, which demonstrates that evaluations by GPT models align closely with human assessments in NLG tasks, we propose utilizing LEAs for interaction evaluation. LEAs assess interactions generated by LEAs and IQAs in Stage 1. The module takes task details, such as questions or articles, and interactions as the input, and output evaluation scores. The prompt contains three parts: (1) role and task description; (2) metrics definition; and (3) evaluation instruction.
@@
Metrics definitions describes the criterion that LEA needs to follow in the evaluation process. They can customized for different tasks. For the question answering task, we add the following prompt to define the “helpfulness” metric:
@@
Finally, all evaluation scores for metrics are calculated by averaging the results of multiple runs. The complete prompt is available in Appendix C.1, and further details about our implementation can be found in Section 4.
@@ 3.3 Assigning the Personas to LEA
Both aforementioned evaluation stages typically use a default persona. While this constitutes a somewhat neutral baseline in knowledge, language proficiency, and beliefs to bo baseline, individual users often exhibit diverse personal preferences and characteristics, making a one-size-fits-all evaluation less effective. Moreover, the persona distribution of the target user group significantly impacts the performance of IQA models in real-world applications. For instance, if 20% of human users prefer brief interactions and 70% prefer detailed information, applying a general LEA to simulate this group of persons is likely to result in poor correlation with downstream users.
@@
For example, when the LEA is assigned with the “Critical-Seeker” persona (definition in C.3) for interaction generation, we adapt the default role and task description (in Stage 1) to:
You prefer interactions rich in critical information. You need help from an assistant and try to get critical information from it to answer the following questions. For interaction evaluation, the default role and task description prompt changes to The AI Assistant should provide straightforward, simple, and concise answers to aid users in deducing solutions. Additionally, the definition of metrics is also adjusted to align with this persona, with further details available in Appendix C.3.
@@ 4.1 Experiment Settings
We apply our evaluation method on the annotated dataset from the study by Lee et al. 29. This dataset consists of 3641 interactions from 331 annotators. Questions in the dataset are multi-choice and are derived from the MMLU dataset 20 (example question in Figure 1). The construction of MMLU requires each worker to participate in 11 random conversations with one of the following three IQA models: TextDavinci (text-davinci-001), TextBabbage (text-babbage-001), and Davinci (davinci-001). At the end of conversations, fluency and helpfulness scores are annotated by annotators. The number of queries and accuracy for each IQA model can be easily deduced from annotations. In this work, We adjust the four metrics to evaluate generated interactions:
@@
To evaluate the effectiveness of IQA-Eval framework, we experiment with different LEA models on the above-mentioned three LLMs IQA models in MMLU. For LEA that conducts both interaction generation and interaction evaluation, we use ChatGPT (GPT-3.5-turbo-1106), GPT4 (GPT-4-1106-preview), Claude (Claude-1).
@@
We report Pearson correlations as the measrue of agreement between the Human evaluations and LEA evaluations of IQA models.
@@ 4.2 Experiment Results
According to Table 2, all models, including GPT4, GPT3.5, and Claude, show high correlation with human evaluations. GPT4 aligns most closely with human judgments in both “Helpfulness” and “Fluency" metrics and the highest overall correlation score. This indicates that these models are capable of effectively performing IQA-Eval framework as LEA models.
@@
In Table 1, ChatGPT scores closest to human judgments, particularly in the "Helpfulness" metric. Conversely, GPT4 and Claude score lower on "Helpfulness" than human evaluations, because they tend to produce inaccurate and repetitive responses that lack coherence and do not directly address user queries, as indicated by their generated explanations. IQA-Eval scores on “Fluency” are close and highly correlated to human judgments. Both scores given by humans and LEA models show that IQA models provide fluent outputs. Furthermore, according to the "# Queries" metric, most models conclude conversations more quickly than humans, except for Claude, which requires more turns, potentially due to its non-OpenAI origins that it needs more turns to adapt the conversational style and understand responses. Notably, GPT4 achieves the highest accuracy among all models. Moreover, we consider the impact of self-enhancement bias and conduct more experiments. Details are in Section 6.4.
@@
Results are in shown Figures 2(a) and 2(b). The Pearson correlation coefficients for fluency and helpfulness between human judgments and LEA evaluations show distinct patterns. ChatGPT demonstrates a stronger correlation with human ratings, recording a correlation score of 0.424 for fluency and a correlation score of 0.306 for helpfulness. In contrast, Claude shows slightly lower correlations (0.281 for fluency and 0.287 for helpfulness). This shows that ChatGPT aligns better with human compared to Claude in these specific metrics.
Moreover, for interactions between humans and IQA models (Figure 2), LEA evaluations moderately correlate with human evaluations. However, for interactions between LEAs and IQA models (Table 2), which is the main focus of our paper, LEA evaluations highly correlate (around 0.6) with human evaluations. This indicates that LEA models are helpful when participating in the whole evaluation process, including both the interaction and evaluation. However, when evaluating interactions between humans and IQA models, LEAs focus differently from humans, which causes moderate correlations.